# Supervised Learning-enhanced Multi-Group Actor Critic for Live Stream Allocation in Feed

### v4

@@ Abstract.
In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to allocate at most one live stream into the video feed for each user request. To maximize long-term user engagement, it is crucial to determine an optimal live stream policy for accurate live stream allocation. The inappropriate live stream allocation policy can significantly affect the duration of the usage app and user retention, which ignores the long-term negative impact of live stream allocation. Recently, reinforcement learning (RL) has been widely applied in recommendation systems to capture long-term user engagement. However, traditional RL algorithms often face divergence and instability problems, which restricts the application and deployment in the large-scale industrial recommendation systems, especially in the aforementioned challenging scenario. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. We also propose a novel reward function to prevent overly greedy live stream allocation. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods under the platform-level constraints but also exhibits enhanced stability in online recommendation scenarios.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live streaming recommendation system, a short video recommendation system, and a blending server. For each user request with timestamp $t$, the live stream recommendation system decides whether to inject the recommended live stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live stream allocation policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration and user retention, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
@@
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in the short video & live stream mixed recommendation system. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
@@
To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B testing to demonstrate its effectiveness and stability.
@@
  • We integrate multi-task supervised learning with traditional critic learning to alleviate TD error accumulation and improve the accuracy of Q-value estimation.
@@ 2. Problem Formulation
  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, online viewer count). We limit the length of user live stream and short video history lists by timestamp order, keeping only the top 50 items.
  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live stream in response to a user’s request. We define the action $a$ as a binary variable, where $a=1$ means injecting a live stream and $a=0$ means not injecting it.
@@
where $\lambda$ is a hyper-parameter to carefully control the final live stream allocation ratio. From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint. Note that we leave the further discussion of the rationality of the above reward function in the Appendix C.
@@ 3.3. Supervised Learning-enhanced Actor Critic
To alleviate the impact of drastic changes in data distribution that could lead to the divergence or instability of RL models, we propose a supervised learning-enhanced actor critic framework to prevent critic networks from getting trapped in model collapse due to large cumulative errors in the critic learning process.
@@ 3.3.1. Layer Normalization
A recent work on RL divergence and instability (Yue et al., 2024) shows that offline critic learning exhibits a self-excitation pattern. Specifically, Q-value iteration can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, which amplifies the TD-error and traps the critic learning process in a positive feedback loop until model collapse. Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate divergence and instability problems.
@@ 3.3.2. Critic Network
where $\hat{Q}^{\prime}(s_{t+1},a_{t+1})=\min_{i=1,2}Q^{\prime}_{\phi_{i}}\left(s_{t+1},a_{t+1}\right)$. We use Huber loss (Hastie et al., 2009) to optimize the above objective.
In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of critic networks. Specifically, due to the instability and divergence issues, target networks often predict inaccurate $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ values, which is much larger than the $r(s_{t},a_{t})$ term. Then the $r(s_{t},a_{t})$ term could not provide any information to guide the critic learning. Therefore, we seamlessly introduce supervised learning (e.g. multi-task learning (Zhang and Yang, 2021)) into the critic learning procedure. Specifically, we divide the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
@@
Formally, we divide the live stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live stream (short video) watch time. Let $y$ denote the actual live stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$. The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$. Then, we have a linear reconstruction layer to reconstruct $y$ from $\delta$:
@@ 4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with a 5000-size embedding layer. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods. Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
@@ 4.7. Online A/B Test Experiment
Note that the Dummy method injects a live stream for each request. We can see that SL-MGAC achieves a significant improvement in live stream watch time while also increasing the app usage duration and user retention than other methods. This shows that SL-MGAC is more effective in maximizing long-term live stream rewards while avoiding the negative impact on app usage duration and user retention. Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may affect the long-term user experience for most users. Moreover, although the dummy method can greatly improve DAU and live stream viewing time, it significantly affects app usage time and user retention.
@@ 5.1. RL in Recommendation Systems
Reinforcement learning (RL) aims to optimize cumulative long-term rewards over time, which has attracted significant attention in the research of recommendation systems in recent years (Afsar et al., 2022). Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend complete item lists, where the number of candidate items can be large. BatchRL-MTF (Zhang et al., 2022), RLUR (Cai et al., 2023), and UNEX-RL (Zhang et al., 2024) leverage RL to model the multi-rank score aggregation process, optimizing the weights for score aggregation. The work in (Miao et al., 2021) explores the use of off-policy RL for personalized search ranking in multiple sessions. CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving the user experience. Furthermore, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
@@ 6. Conclusion
In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream allocation policies under the platform-level constraints on app usage duration and user retention in the challenging context of a live stream & short video mixed recommendation system. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation. Compared with existing RL methods for recommendation systems, our approach not only focuses on improving online performance, but also improves the stability and robustness of the RL model. In practice, we minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully implemented in large-scale industrial recommendation systems. Currently, we are trying to apply the SL-MGAC methods to other similar recommendation scenarios, including E-commerce and advertising. We also aim to continuously improve the performance of SL-MGAC in further research. Learning live stream policy from a long sequence of user behaviors $(s_{t},a_{t},r_{t})$ and exploring more refined and complex user state representations would lead to our future work.
7. Acknowledgment
The authors acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.

### v5

@@ Abstract.
In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to allocate at most one live stream to the video feed for each user request. The inappropriate policy which ignores the long-term negative impact of live stream allocation can significantly affect app usage duration and user retention. To maximize long-term user engagement, it is crucial to determine an optimal policy for accurate live stream allocation. Recently, reinforcement learning (RL) has been widely applied in recommendation systems to capture long-term user engagement. However, traditional RL algorithms often face divergence and instability problems, which restricts application and deployment in large-scale industrial recommendation systems, especially in the aforementioned challenging scenario. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task supervised reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. We also propose a novel reward function to prevent overly greedy live stream allocation. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods under platform-level constraints, but also exhibits improved stability in online recommendation scenarios.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2, August 3–7, 2025, Toronto, ON, Canada; Toronto, ON; Canadaisbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live streaming recommendation system, a short video recommendation system, and a blending server. For each user request with timestamp $t$, the live stream recommendation system decides whether to inject the recommended live stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live stream allocation policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on app usage duration and user retention, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
@@
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in the short video & live stream mixed recommendation system. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023). Most importantly, the live-stream allocation agent is the final module of the live-stream RS, as shown in Fig 1. If the RL agent is very instable or collapse in the online environment, we may inject too many live-streams in the short-video feed, which will cause severe malfunctions and will highly affect the user retention. Hence, in the industrial RS, this RL application is more risky than existing RL methods (Cai et al., 2023; Zhang et al., 2022, 2024), which are only applied in the multi-rank score aggregation of the ranking stage.
@@
To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B tests to demonstrate its effectiveness and stability.
@@
  • We integrate multi-task supervised reward learning with traditional critic learning to alleviate TD error accumulation and improve the accuracy of Q-value estimation.
@@ 2. Problem Formulation
  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, etc). We limit the length of user live stream and short video history lists by timestamp order, keeping only the top 50 items.
  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live stream in response to a user’s request. We define the action $a$ as a binary variable, where $a=1$ means injecting a live stream.
@@
where $\lambda$ is a hyper-parameter to carefully control the final live stream allocation ratio. From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint. Note that we leave the further discussion of the above reward function in the Appendix C.
@@ 3.3. Supervised Learning-enhanced Actor Critic
To alleviate the impact of drastic changes in the data distribution that could lead to divergence or instability of RL models, we propose a supervised learning-enhanced actor critic framework to prevent critic networks from being trapped in model collapse due to large cumulative errors in the critic learning process.
@@ 3.3.1. Layer Normalization
A recent work on RL divergence and instability (Yue et al., 2024) shows that offline critic learning exhibits a self-excitation pattern. Specifically, iteration of the Q value can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, amplifying the TD error and trapping the critic learning process in a positive feedback loop until the model collapse. Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate divergence and instability problems.
@@ 3.3.2. Critic Network
where $\hat{Q}^{\prime}(s_{t+1},a_{t+1})=\min_{i=1,2}Q^{\prime}_{\phi_{i}}\left(s_{t+1},a_{t+1}\right)$. We use the Huber loss (Hastie et al., 2009) to optimize the above objective.
In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of critic networks. Specifically, due to instability and divergence issues, target networks often predict inaccurate $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ values, which is much larger than the $r(s_{t},a_{t})$ term. Then the term $r(s_{t},a_{t})$ could not provide any information to guide the learning of the critic. Therefore, we seamlessly introduce supervised learning (e.g. multi-task learning (Zhang and Yang, 2021)) into the critic learning procedure. Specifically, we divide the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
@@
Formally, we divide the live stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live stream (short video) watch time. Let $y$ denote the actual live stream or short video watch time, which falls in the time bin $[y_{st},y_{end}]$. The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$. Then, we have a linear reconstruction layer to reconstruct $y$ from $\delta$:
@@ 4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with a 5000-size embedding layer. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods and $\lambda$ in Eq 5 is 0.1. Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
@@ 4.7. Online A/B Test Experiment
Note that the Dummy method injects a live stream for each request. We can see that SL-MGAC achieves a significant improvement in live stream watch time while also increasing the app usage duration and user retention than other methods. This shows that SL-MGAC is more effective in maximizing long-term live stream rewards while avoiding the negative impact on app usage duration and user retention. Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may affect the long-term user experience for most users. Moreover, although the dummy method can greatly improve DAU and live stream viewing time, it significantly affects app usage duration and user retention.
@@ 5.1. RL in Recommendation Systems
Reinforcement learning (RL) aims to optimize cumulative long-term rewards over time, which has attracted significant attention in the research of recommendation systems in recent years (Afsar et al., 2022). Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend complete item lists, where the number of candidate items can be large. BatchRL-MTF (Zhang et al., 2022), RLUR (Cai et al., 2023), and UNEX-RL (Zhang et al., 2024) leverage RL to model the multi-rank score aggregation process, optimizing the weights for score aggregation. The work in (Miao et al., 2021) explores the use of off-policy RL for personalized search ranking in multiple sessions. CrossDQN (Liao et al., 2022) introduces an RL-based approach to ad allocation in a feed, aiming to maximize revenue while improving the user experience. Furthermore, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
@@ 6. Conclusion
In the challenging context of a live stream & short video mixed recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream allocation policies under the platform-level constraints on app usage duration and user retention. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task supervised reward learning with traditional critic learning to constrain Q-value estimation. Compared with existing RL methods for recommendation systems, our approach not only focuses on improving online performance, but also improves the stability and robustness of the RL model. In practice, we minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully implemented in large-scale industrial recommendation systems. Currently, we are trying to apply the SL-MGAC methods to other similar recommendation scenarios, including E-commerce and advertising. We also aim to continuously improve the performance of SL-MGAC in further research. Learning live stream policy from a long sequence of user behaviors $(s_{t},a_{t},r_{t})$ and exploring more refined and complex user state representations would lead to our future work.
@@ Appendix C More Ablation Studies
We conduct further online AB experiments to demonstrate the performance of the SL-MGAC policy at different values of $\lambda$ in Eq 5, as shown in Table 5. It indicates that when $\lambda>0$, the live stream allocation policy of SL-MAGC behaves more conservatively to balance live stream watch time and video watch time, and finally optimize user’s long-term engagement and retention. However, when $\lambda<0$, the SL-MAGC behaves more greedily to allocate live streams to the video feed. Specifically, we find that the SL-MAGC allocates more live streams to the requests that occur earlier in a session due to the position bias on the reward, which would interrupt the user interest throughout the session and hence affect user retention.
$\lambda$ live stream DAU live stream Watch Time Video Watch Time App Usage Duration User Retention
0.2 -1.245% -3.567% +0.215% +0.123% +0.106%
0.1 +2.616% +7.431% +0.197% +0.121% +0.086%
-0.1 +4.123% +6.564% -0.248% -0.362% -0.112%
-0.2 +5.876% +7.215% -0.296% -0.445% -0.159%
Table 5. Online AB Test Performance of SL-MGAC under different values of $\lambda$.